How to derive 3 equations of motion using calculus??????

Grade:11

1 Answers

Naveen sharma
5 years ago
V=u+at,s=ut+1/2at²andv²=u²+2asThis is the equation ofCalculus method Hu j De xjnxjxnxnmxnxnj njdjrifiiaoaoyfyduoppsjhdkdl. Bcbfmskxdbjdlc